OpenStudy (anonymous):

Could someone help me with finding the ratios for sin x and cos x in a right triangle where the hypotenuse is 12, the opposite is 5, and the adjacent is the square root of 119 (10.9 if I round/simplify it, but...)?

OpenStudy (anonymous):

sinx = opposite/hypotenus = 5/12 cosx = adjacent/hypotenuse = sqrt(119)/12

OpenStudy (anonymous):

Thank you! c: